OpenStudy (anonymous):

how does one change the integration limit from 0 to infinitiy to -1 and +1

OpenStudy (agent0smith):

Not sure what you mean... you'd change them by.. changing them.

OpenStudy (anonymous):

I have an equation where the integration is for the radial part but i need it to fit the limits to cos(theta)

OpenStudy (agent0smith):

Maybe just post the question...

OpenStudy (anonymous):

\[\int\limits_{o}^{\inf}\frac{ V(r)rP(\cos \theta) }{ 2r _{1}r _{2} }dr\]

OpenStudy (anonymous):



OpenStudy (anonymous):

this is the diagram that i consider to make a substitution